OpenStudy (anonymous):

evaluate int frac{dx}{4x sqrt{x ^{2}}-4^{2}}

OpenStudy (anonymous):

\[evaluate \int\limits \frac{dx}{4x \sqrt{x ^{2}}-4^{2}}\]

OpenStudy (lgbasallote):

isnt the square rootof x^2 just x \[\int\frac{dx}{4x^2 - 4^2}\] rewrite \[\int \frac{dx}{(2x)^2 - 4^2}\] does that look familiar?

OpenStudy (anonymous):

i think its \[\int\limits \frac{dx}{4x \sqrt{x^2-16}} \]